OpenStudy (anonymous):

Where is the vertex of the parabola? y = x2 + 2x + 3 A. Above the x-axis B. Below the x-axis C. On the x-axis D. On the y-axis

OpenStudy (anonymous):

y= x^2 + 2x +3 is what I meant.

OpenStudy (aravindg):

First of all find the discriminant

OpenStudy (aravindg):

Did you find it ?

OpenStudy (anonymous):

no?

OpenStudy (aravindg):

Do you know how to find discriminant?

OpenStudy (anonymous):

No.

OpenStudy (aravindg):

For a quadratic ax^2+bx+c Discriminant=b^2-4ac

OpenStudy (anonymous):

Now what?

OpenStudy (aravindg):

Found it ?

OpenStudy (anonymous):

Omg, I'm so confused! -.-

OpenStudy (aravindg):

Why ?

OpenStudy (anonymous):

Nvm -_____-

OpenStudy (aravindg):

Take an example .For quadratic x2+4x+5 Discriminant=16-4*1*5=1 6-20=-4

OpenStudy (anonymous):

Okayyyy? So what's the answer?

OpenStudy (aravindg):

Use that in your question ...I will verify